OpenStudy (anonymous):

Is the equation true, false, or open? 4y+8=6y+3 a. true the expressions are the same for all values of the variables. b. false the expressions are never the same c. open there is a variable

OpenStudy (igreen):

4y + 8 = 6y + 3 Subtract 8 to both sides: 4y = 6y - 5 Subtract 6y to both sides: -2y = -5 Divide -2 to both sides: y = -2/5 or 0.4

OpenStudy (igreen):

Sorry that's -5/-2 or 5/2 or 2.5

OpenStudy (igreen):

So your answer's going to be Open, because there is a variable and a solution.

OpenStudy (anonymous):

which value is a solution of the equation 3-2x=-5

OpenStudy (anonymous):

@iGreen